### v3

# On Computing Minimum Selective Subsets in Unit Disk and Circle Graphs

Abstract
In a connected simple graph $G=(V(G),E(G))$, each vertex is assigned one of $c$ colors, where $V(G)=\bigcup_{\ell=1}^{c}V_{\ell}$ and $V_{\ell}$ denotes the set of vertices of color $\ell$. A subset $S\subseteq V(G)$ is called a selective subset if, for every $\ell$, $1\leq\ell\leq c$, every vertex $v\in V_{\ell}$ has at least one nearest neighbor in $S\cup(V(G)\setminus V_{\ell})$ that also lies in $V_{\ell}$. The Minimum Selective Subset (MSS) problem asks for a selective subset of minimum size.
We show that the MSS problem is log-APX-hard on general graphs, even when $c=2$. As a consequence, the problem does not admit a polynomial-time approximation scheme (PTAS) unless $\textup{{P}}=\textup{{NP}}$. On the positive side, we present a PTAS for unit disk graphs that does not require a geometric representation and applies for arbitrary $c$. We further prove that MSS remains NP-complete in unit disk graphs for arbitrary $c$. In addition, we show that the MSS problem is APX-hard on circle graphs, even when $c=2$  [^1] This work appeared in CALDAM 2026 [21].
1 Introduction
Many computational tools have been developed for supervised learning methods on a labeled training set $T$ embedded in a metric space $(X,d)$. Each data point $t\in T$ is associated with a label (also called a character or color), chosen from a set $C=\{1,2,\dots,c\}$. The objective is to extract a smallest possible subset $S\subseteq T$ such that every point in $T$ either belongs to $S$ or has at least one nearest neighbor (with respect to the metric $d$) within $S$ that shares the same character. This optimization problem, called the Minimum Consistent Subset (MCS), was originally formulated by Hart [16] in 1968, which has received thousands of citations, highlighting its significant impact in the field. However, the paper [16] did not establish any complexity results or algorithms.
Later in 1991, Wilfong [27] defined two problems MCS and MSS together and proved that the MCS and MSS problems are NP-complete in $\mathbb{R}^{2}$ for $c\geq 3$ and $c\geq 2$, respectively. It also proposed a polynomial-time algorithm when there is only one red point and all other points are blue in $\mathbb{R}^{2}$. Later, in 2018, it was proved that MCS remains NP-complete when $c=2$ in $\mathbb{R}^{2}$ [17]. Recently, Banerjee et al. [2] showed that MCS is W[2]-hard (for arbitrary $c$) and MSS is W[1] (for $c=2$), both parameterized by the solution size. Various algorithms, including those for many restricted inputs for the MCS problem in $\mathbb{R}^{2}$ have been proposed [2, 5, 7], highlighting its significance in machine learning and computational geometry. The only algorithm for the MSS problem is a PTAS, which was established when $c=2$ [2].
The Minimum Selective Subset (MSS) problem plays a crucial role in optimizing data selection by identifying the smallest subset which preserves essential information. It can be viewed both as a clustering and a proximity problem. This is particularly useful in applications such as fingerprint recognition, character recognition, and pattern recognition, where it helps reduce redundancy and improve decision-making in classification and feature selection tasks. MSS was introduced because the standard MCS methods, like Hart’s [16], could not guarantee that the resulting subset was the smallest possible size. MSS applies a stronger consistency condition on the subset to create a mathematical framework that is amenable to an exact minimization procedure, thereby fulfilling the optimization goal that MCS heuristics failed to satisfy. So far, we have discussed MCS and MSS problems along with their published results in $\mathbb{R}^{2}$. We now turn to these problems in the context of graph algorithms.
Banerjee et al. [2] proved that MCS is W[2]-hard [9] when parameterized by the solution size, even with only two colors on general graphs. Dey et al. [11, 12] provided polynomial-time algorithms for MCS on some simple graph classes including paths, spiders, caterpillars, combs, and trees (for trees, $c=2$). XP, NP-complete, and FPT (when $c$ is a parameter) results on trees, can be found in [1, 4]. The MCS problem is also NP-complete on interval graphs [4] and APX-hard on circle graphs [18]. Variants, such as the Minimum Consistent Spanning Subset (MCSS) and the Minimum Strict Consistent Subset (MSCS) of MCS, have been studied on trees [3, 6, 19]. However, the algorithmic results for MSS have not been extensively studied to date. Banerjee et al. [2] only showed that MSS is NP-complete on general graphs. Very recently, the MSS problem has been studied in various settings, including $\mathcal{O}(\log n)$-approximation algorithms for general graphs, NP-complete results for planar graphs, and linear-time algorithms for trees and unit interval graphs [20] (published in CCCG 2025).
Our Contributions. The MSS problem admits an $\mathcal{O}(\log n)$-approximation on general graphs, which raises the question of whether better approximations exist. We show in Section 3 that MSS is log-APX-hard even when $c=2$. Hence, the problem is also APX-hard and does not admit a PTAS on general graphs. This leads to the natural question of whether some graph classes allow a PTAS. To date, none are known. We answer this by proving in Section 5 that MSS admits a PTAS on unit disk graphs for arbitrary $c$, without requiring a geometric representation. Unit disk graphs are also fundamental in wireless networks, robotics, and computational geometry, where efficient approximation algorithms are highly relevant [14]. Before presenting our PTAS, we establish in Section 4 that MSS remains NP-complete on unit disk graphs when $c$ is arbitrary. We also investigate whether MSS is APX-hard in other graph classes. In Section 6, we prove that MSS is APX-hard on circle graphs even when $c=2$. Circle graphs, which model intersecting chords, have applications in VLSI design, scheduling, and bioinformatics [26, 24]. All proofs of results marked with $(^{\ast})$ can be found in the Appendix.
2 Preliminaries
Let $G=(V(G),E(G))$ be a graph, where $V(G)$ is the vertex set and $E(G)$ is the edge set. For any $U\subseteq V(G)$, $G[U]$ denotes the subgraph of $G$ induced on $U$, and $\lvert U\rvert$ is the cardinality of $U$. We denote $[n]$ as the set of integers $\{1,\ldots,n\}$. We use an arbitrary vertex color function $C:V(G)\rightarrow[c]$, which assigns each vertex exactly one color from the set $[c]$. For a subset of vertices $U\subseteq V(G)$, let $C(U)$ represent the set of colors of the vertices in $U$, formally defined as $C(U)=\{C(u)\mid u\in U\}$. The shortest path distance (i.e., hop-distance) between two vertices $u$ and $v$ in $G$ is denoted by $\mathrm{d}(u,v)$. Distance between $v\in V(G)$ and the set $U\subseteq V(G)$ is given by $\mathrm{d}(v,U)=\min_{u\in U}\mathrm{d}(v,u)$. Similarly, the distance between two subgraphs $G_{1}$ and $G_{2}$ in $G$ is defined as $\mathrm{d}(G_{1},G_{2})=\min\{\mathrm{d}(v_{1},v_{2})\mid v_{1}\in V(G_{1}),v_{2}\in V(G_{2})\}$. The set of nearest neighbors of $v$ in the set $U$ is denoted as $\mathrm{\hat{N}}(v,U)$, formally defined as $\mathrm{\hat{N}}(v,U)=\{u\in U\mid\mathrm{d}(v,u)=\mathrm{d}(v,U)\}$. Therefore, if $v\in U$, then $\mathrm{\hat{N}}(v,U)=\{v\}$. The set of vertices in $U$ adjacent to $v$ is given by $\mathrm{N}(v,U)=\{u\in U\mid(u,v)\in E(G)\}.$ We also define $\mathrm{N}[v,U]=\{v\}\cup\mathrm{N}(v,U).$ For any two subsets $U_{1},U_{2}\subseteq V(G)$, we define $\mathrm{N}(U_{1},U_{2})=\bigcup_{v\in U_{1}}\mathrm{N}(v,U_{2}),$ and $\mathrm{N}[U_{1},U_{2}]=\bigcup_{v\in U_{1}}\mathrm{N}[v,U_{2}].$ Most symbols and notations follow standard conventions from [13]. Suppose $G=(V(G),E(G))$ is a given simple connected undirected graph where $\bigcup_{i=1}^{c}V_{i}=V(G)$ and $V_{i}\cap V_{j}=\emptyset$ for $i\neq j$ and each vertex in $V_{i}$ is assigned color $i$. A Minimum Consistent Subset (MCS) is a subset $S\subseteq V(G)$ of minimum cardinality such that for every vertex $v\in V(G)$, if $v\in V_{i}$, then $\mathrm{\hat{N}}(v,S)\cap V_{i}$ is non-empty.
Definition 1(Selective Subset)
A subset $S\subseteq V(G)$ is called a Selective Subset (MSS) if, for each vertex $v\in V(G)$, if $v\in V_{i}$, the set of nearest neighbors of $v$ in $S\cup(V(G)\setminus V_{i})$, denoted as $\mathrm{\hat{N}}(v,S\cup(V(G)\setminus V_{i}))$, contains at least one vertex $u$ such that $C(v)=C(u)$. An MSS is a selective subset of minimum cardinality. The decision version of the MSS problem is as follows:
Decision Version of Selective Subset Problem on Graphs Input: A graph $G=(V(G),E(G))$, a coloring function $C:V(G)\rightarrow[c]$, and an integer $s$. Question: Does there exist a selective subset of size at most $s$ for $(G,C)$?
In other words, we seek a vertex set $S\subseteq V(G)$ of minimum cardinality such that every vertex $v$ has at least one nearest neighbor of the same color in the graph, excluding vertices of the same color as $v$ that are not in $S$. If all the vertices of a graph $G$ are of the same color (i.e., $G$ is monochromatic), then any vertex in the graph forms a valid MSS. Figure 1 illustrates an example of MSS.
Figure 1: Colors: blue $\equiv$ cross, green $\equiv$ disk, red $\equiv$ square, and orange $\equiv$ circle. $V(G)=$ $V_{blue}$ $\cup$ $V_{green}$ $\cup$ $V_{red}\cup V_{orange}$, where $V_{blue}=$ $\{v_{1}\}$, $V_{green}$ == $\{v_{2}$ $\dots,$ $v_{9},v_{23},$ $\dots,v_{26}\}$, $V_{red}=\{v_{10},\dots,v_{15},v_{27},\dots,v_{30}\}$, and $V_{orange}=\{v_{16},\dots,v_{22},v_{31},v_{32}\}$. $S=$ $\{v_{1},$ $v_{2},$ $v_{3},$ $v_{7},$ $v_{8},$ $v_{9},v_{10},v_{11},v_{15},v_{16}\}$ is an MSS, and $S=\{v_{1},$$v_{4},$ $v_{5},$ $v_{7},v_{8},v_{9},v_{12},v_{13},v_{15},v_{16}\}$ is also an MSS. Brown-dotted regions indicate the blocks. The complete list of blocks is $B_{1}=\{v_{1}\}$, $B_{2}=\{v_{2},\dots,v_{7},v_{23},\dots,v_{26}\}$, $B_{3}=\{v_{10},\dots,v_{15},v_{27},\dots,v_{30}\}$, $B_{4}=\{v_{16},\dots,v_{22},v_{31},v_{32}\}$, $B_{5}=\{v_{8}\}$, $B_{6}=\{v_{9}\}$. $B_{2,1}=\{v_{2},v_{3},v_{6},v_{7}\}$, $B_{2,2}=\{v_{4},v_{5},v_{25},v_{26}\}$. $\{\{v_{2}\},\{v_{3}\},\{v_{7}\}\}$ is a collection of 2-distance sets in $B_{2,1}$.
Definition 2(Block)
A block is a maximal connected subgraph whose vertices all have the same color (i.e., a maximal connected monochromatic subgraph).
Figure 1 illustrates an example of the blocks. Suppose $B_{1},\dots,B_{k}$ is the complete list of blocks in $G$. We assume that $\lvert V(G)\rvert=n$, so that $k\leq n$. We form the sets $B_{i}^{1}$, $B_{i}^{2}$, $B_{i,3}$ for each $i=1,\dots,k$ as follows (see Figure 1):
  • Initially, $B_{i,1}:=\emptyset$, $B_{i,2}:=\emptyset$.
  • For each vertex $v\in B_{i}$, if there exists a vertex $u\in\mathrm{N}(v,V(G))$ such that $C(u)\neq C(v)$, then $v\in B_{i,1}$.
  • For any vertex $v\in B_{i}\setminus B_{i,1}$ if $\mathrm{d}(v,B_{i,1})=1$, then $v\in B_{i,2}$.
  • We denote $B_{i,3}=B_{i,1}\cup B_{i,2}$.
Lemma 1
For any vertex $v\in B_{i,1}$ and a selective subset $S$, we have $\mathrm{N}[v,B_{i,3}]\cap S\neq\emptyset$ for $1\leq i\leq k$.
3 log-APX-hardness of MSS on General Graphs
We establish a reduction from the Minimum Dominating Set (MDS) problem to the MSS problem. In the Minimum Dominating Set problem, the input is a graph $G$ together with an integer $s$. The task is to decide whether there exists a subset $D\subseteq V(G)$ of size at most $s$ such that every vertex $u\in V(G)$ satisfies $\mathrm{N}[u,V(G)]\cap D\neq\emptyset$. It is well known that the Minimum Set Cover problem is log-APX-hard (see [25] for complexity class definitions), and moreover, it is NP-hard to approximate it within a factor of $\delta\cdot\log n$ for some positive constant $\delta$ [23]. Since there exists an $L$-reduction from the Minimum Set Cover problem to the Minimum Dominating Set problem, the Minimum Dominating Set problem is also log-APX-hard.
Let $(G,s)$ be an arbitrary instance of the Minimum Dominating Set problem. We construct an instance $(G^{\prime},C,s+1)$ for the MSS problem as follows (see Figure 2(a)). Define the new graph $G^{\prime}$ with $V(G^{\prime})=V(G)\cup\{z\}$ and $E(G^{\prime})=E(G)\cup\{(z,u)\mid u\in V(G)\}$. The color function $C$ assigns color 1 to all vertices $u\in V(G)$, and color 2 to the additional vertex $z$.
Lemma 2
$G$ has a dominating set of size at most $s$ if and only if $G^{\prime}$ has a selective subset of size at most $s+1$.
Figure 2: Colors: blue $\equiv$ disk, red $\equiv$ square, and green $\equiv$ circle. (a) Reduction from an instance of Minimum Dominating Set problem to an instance of MSS problem when $c=2$. (b) Example of the reduction when $c=3$.
Theorem 3.1
There exists a constant $\delta>0$ such that it is NP-hard to approximate the MSS problem within a factor of $\delta\cdot\log n$, where $n$ is the number of vertices in the graph.
Remark 1
Note that the above reduction remains valid even if we add any number of new vertices (each adjacent to all of $V(G)$) and assign each a distinct color. The correctness of Lemma 2 and the resulting hardness theorem continue to hold under this extended construction (see Figure 2(b)).
4 NP-completeness of MSS on Unit Disk Graphs
A graph $U=(V(U),E(U))$ is called a unit disk graph (UDG) if its vertices can be represented as points in the Euclidean plane such that an edge exists between two vertices if and only if their Euclidean distance is at most $2$.
Formally, $U$ is a UDG if there exists a mapping $f:V\to\mathbb{R}^{2}$ such that:
$(u,v)\in E(U)\iff\|f(u)-f(v)\|\leq 2$ (1)
where $\|\cdot\|$ denotes the Euclidean norm.
Clark et al. [8] showed that Minimum Dominating Set (MDS) is NP-complete in UDG. We reduce from an instance $U$ of the UDG with $\lvert V(U)\rvert=n$ to an instance $U^{\prime}$ as follows:
Figure 3: Colors: blue $\equiv$ disk, red $\equiv$ circle, orange $\equiv$ fsquare, green $\equiv$ dash dotted circle, brown $\equiv$ dash dotted fsquare, darkred $\equiv$ square, darkgreen $\equiv$ cross, lightgreen $\equiv$ dotted circle, and violet $\equiv$ dotted square. An example of the reduction when $m=1$. Each blue disk in $U^{\prime}$ is adjacent to a disk of a distinct color, different from all other colors in $U^{\prime}$.
Reduction. Define $V(U^{\prime})=V(U)\cup V(X)$ and $E(U^{\prime})=E(U)\cup E(X)$, where we introduce a subgraph $X$ with vertex set $V(X)$ and edge set $E(X)$ as follows (see Figure 3). Initially, set $V(X):=\emptyset$ and $E(X):=\emptyset$. Assign color 0 to all unit disks in $V(U)$, i.e., $C(V(U))=\{0\}$. For each $u_{i}\in V(U)$, introduce a total of $m$ (where $m\in\mathbb{N}$) unit disks $v_{i}^{1},\dots,v_{i}^{m}\in V(X)$. Additionally, place $v_{i}^{1},\dots,v_{i}^{m}$ in such a way that $u_{i}$ and $v_{i}^{l}$ are adjacent for $1\leq l\leq m$ (where $v_{i}^{l}$ may also be adjacent to other unit disks). These new edges are added to $E(X)$. All unit disks in $V(X)$ have distinct colors, none of which is 0. Thus, we obtain $\lvert V(X)\rvert=mn$, $\lvert V(U^{\prime})\rvert=nm+n$, and $\lvert C(V(U^{\prime}))\rvert=nm+1$.
Lemma 3
$U$ has a dominating set of size $t$ if and only if $U^{\prime}$ has a selective subset of size $nm+t$.
Theorem 4.1
MSS is NP-complete on unit disk graphs.
5 PTAS of MSS on Unit Disk Graphs
A unit disk graph may admit multiple geometric representations. In this work, we assume that the geometric representation $f$ is either unknown or not explicitly given. We first describe our approach for a general graph before restricting to unit disk graphs. If $G$ is not a connected graph, we apply the algorithm independently on each component; hence, we may assume that $G$ is connected.
The key idea is that blocks are independent within any selective subset solution. Exploiting this property, we compute local selective subsets within each block independently. For the set $B_{i,1}$ corresponding to a given block $B_{i}$, we define a family of sets $D_{i,1}^{1},\dots,D_{i,1}^{t_{i}}$ such that $\mathrm{d}(D_{i,1}^{j},D_{i,1}^{\ell})>2$ for all $j\neq\ell$. By Lemmas 1 and 4, this ensures that no two sets $D_{i,1}^{j}$ and $D_{i,1}^{\ell}$ share a common vertex in optimal solution. Consequently, the solutions for $D_{i,1}^{1},\dots,D_{i,1}^{t_{i}}$ provide a lower bound on the size of the optimal solution. However, the union of these solutions does not necessarily form a valid solution for the block $B_{i}$. To address this, we enlarge each set $D_{i,1}^{j}$ into a corresponding set $E_{i,1}^{j}$, ensuring that the union of the solutions for $E_{i,1}^{1},\dots,E_{i,1}^{t_{i}}$ yields a valid solution for $B_{i}$. We refer to the solutions for $D_{i,1}^{j}$ and $E_{i,1}^{j}$ as local solutions. By combining these local solutions, we obtain a blockwise selective subset, and taking the union over all blocks yields a global solution used in our PTAS.
For any subset $U\subseteq V(G)$, we denote its minimum selective subset (local) by $S^{\min}(U)$ and a selective subset (local) of $U$ by $S(U)$. The global optimum is denoted by $S^{\min}$.
Lemma 4
For any minimum selective subset $S^{min}$ of $G$, we have
  • No vertex $v\in B_{i}\setminus B_{i,3}$ belongs to $S^{min}$.
  • $S^{min}\subseteq\bigcup_{i=1}^{k}B_{i,3}$.
Lemma 1 ensures that for each vertex $v\in B_{i,1}$, either $v\in S^{min}$ or at least one of its adjacent vertices in $B_{i,3}$ belongs to $S^{min}$. Importantly, the choice of including $v$ itself or one of its adjacent vertices from $B_{i,3}$ in $S^{min}$ does not affect the selection of vertices in other blocks. Combined with Lemma 4, which establishes that $S^{\min}\subseteq\bigcup_{i=1}^{k}B_{i,3}$, we obtain the following remark:
Remark 2
The blocks are independent in constructing a selective subset; that is, the selection of vertices in one block does not constrain the selection in other blocks.
We now apply an appropriate algorithm to each block separately due to Remark 2. To do this, we first define a selective subset for each block as follows.
Definition 3(Selective Subset of $B_{i,1}$)
A selective subset of $B_{i,1}$, denoted by $S(B_{i,1})$, is a subset of $B_{i,3}$ such that for every vertex $v\in B_{i,1}$, either $v\in S(B_{i,1})$ or $\mathrm{N}(v,B_{i,3})\cap S(B_{i,1})\neq\emptyset$. By Lemmas 1 and 4, it follows that $S(B_{i,1})\subseteq\mathrm{N}[B_{i,1},B_{i,3}]\subseteq B_{i,3}$.
Theorem 5.1
Let $G$ be a connected graph with blocks $B_{1},\dots,B_{k}$, and for each block $B_{i}$ let $S(B_{i,1})$ be any selective subset of $B_{i,1}$. Let $S=\bigcup_{i=1}^{k}S(B_{i,1})$, then $S$ is a selective subset of $G$. Moreover, if each $S(B_{i,1})$ satisfies $|S(B_{i,1})|\leq(1+\epsilon)\,|S^{\min}(B_{i,1})|,$ then $|S|\leq(1+\epsilon)\,|S^{\min}|$, where $S^{\min}$ denotes a minimum selective subset of $G$.
5.1 Finding Local Selective Subsets
We now establish a bound on a local solution using 2-distance subsets for our problem and then merge all local solutions to obtain the desired solution.
Definition 4(2-distance Subsets)
A collection of subsets of the vertices in $B_{i,1}$, denoted as $D_{i}=\{D_{i,1}^{1},\dots,D_{i,1}^{t_{i}}\}$, is called a collection of 2-distance subsets if the following properties hold (see example in Figure 1):
  • $D_{i,1}^{j}\subseteq B_{i,1}$ for all $1\leq j\leq t_{i}$.
  • The subgraph $G[D_{i,1}^{j}]$ is connected in the induced subgraph $G[B_{i,3}]$, i.e., the induced subgraph $G[D_{i,1}^{j}]$ may not be connected itself, but any two vertices in $D_{i,1}^{j}$ must have a path between them in $G[B_{i,3}]$.
  • The subsets are pairwise at a distance greater than two in $G[B_{i,3}]$, i.e., $\mathrm{d}(D_{i,1}^{j},D_{i,1}^{l})>2$ in the subgraph $G[B_{i,3}]$ when $j\neq l$.
Definition 5(Local Selective Subset)
A local selective subset of $D_{i,1}^{j}$, denoted by $S(D_{i,1}^{j})$, is a subset of $B_{i,3}$ such that for every vertex $v\in D_{i,1}^{j}$, either $v\in S(D_{i,1}^{j})$ or $\mathrm{N}(v,B_{i,3})\cap S(D_{i,1}^{j})\neq\emptyset$. By Lemmas 1 and 4, it follows that $S(D_{i,1}^{j})\subseteq\mathrm{N}[D_{i,1}^{j},B_{i,3}]\subseteq B_{i,3}$.
Lemma 5
For any $j\neq l$, the following holds:
  • $\mathrm{N}[D_{i,1}^{j},B_{i,3}]\cap\mathrm{N}[D_{i,1}^{l},B_{i,3}]=\emptyset$.
  • $S^{min}(D_{i,1}^{j})\cap S^{min}(D_{i,1}^{l})=\emptyset$.
  • $\big(S^{min}\cap S^{min}(D_{i,1}^{j})\big)\cap\big(S^{min}\cap S^{min}(D_{i,1}^{l})\big)=\emptyset.$
Lemma 5 implies that the solutions $S^{min}(D_{i,1}^{j})$ and $S^{min}(D_{i,1}^{l})$ do not share a common vertex in $S^{min}$ for $j\neq l$.
Lemma 6
$S^{min}\cap\mathrm{N}[D_{i,1}^{j},B_{i,3}]$ is a local selective subset of $D_{i,1}^{j}$.
Lemma 7
For any collection of 2-distance subsets $D_{i}=\{D_{i,1}^{1},$$\dots,$ $D_{i,1}^{t_{i}}\}$, where $1\leq i\leq k$ in the graph $G$; we have: $\sum_{i=1}^{k}\sum_{j=1}^{t_{i}}\lvert S^{min}(D_{i,1}^{j})\rvert\leq\lvert S^{min}\rvert$.
Lemma 7 shows that 2-distance subsets yield a lower bound on the size of a minimum selective subset. However, the set $\sum_{i=1}^{k}\sum_{j=1}^{t_{i}}S^{min}(D_{i,1}^{j})$ need not form a selective subset of the entire graph $G$. To construct a selective subset for $G$, we enlarge each $D_{i,1}^{j}$ to a corresponding set $E_{i,1}^{j}$ that remains locally bounded while still providing a valid local solution. This enlargement allows us to approximate a selective subset of $G$.
Theorem 5.2
Let $D_{i}=\{D_{i,1}^{1},\dots,D_{i,1}^{t_{i}}\}$ be a collection of 2-distance subsets, and $\{E_{i,1}^{1},\dots,E_{i,1}^{t_{i}}\}$ be the corresponding collection of subsets of $B_{i,1}$ such that $D_{i,1}^{j}\subseteq E_{i,1}^{j}$ for all $1\leq i\leq k$ and $1\leq j\leq t_{i}$.
If there exists a bound $\delta\geq 1$ such that $\lvert S^{min}(E_{i,1}^{j})\rvert\leq\delta\cdot\lvert S^{min}(D_{i,1}^{j})\rvert$ for all $1\leq i\leq k$ and $1\leq j\leq t_{i}$, and if $\bigcup_{i=1}^{k}\bigcup_{j=1}^{t_{i}}S^{min}(E_{i,1}^{j})$ forms a selective subset of $G$, then $\bigcup_{i=1}^{k}\bigcup_{j=1}^{t_{i}}S^{min}(E_{i,1}^{j})$ is a $\delta$-approximation of a minimum selective subset of $G$.
5.2 Finding a Global Selective Subset
For $r=0,1,\dots$, we recursively define the $r$-th neighborhood of any vertex $v\in B_{i,1}$ in $B_{i,3}$ by
$\mathrm{N}_{i}^{r}[v,B_{i,3}]=\mathrm{N}[\mathrm{N}_{i}^{r-1}[v,B_{i,3}],B_{i,3}],$
with
$\mathrm{N}_{i}^{0}[v,B_{i,3}]=\{v\},\quad\mathrm{N}_{i}^{1}[v,B_{i,3}]=\mathrm{N}[v,B_{i,3}].$
Since $\mathrm{N}_{i}^{r}[v,B_{i,3}]\subseteq B_{i,3}$, we partition $\mathrm{N}_{i}^{r}[v,B_{i,3}]$ into $X_{i}^{r}\subseteq B_{i,1}$ and $Y_{i}^{r}\subseteq B_{i,2}$. We will later use $X_{i}^{r}$ and $Y_{i}^{r}$ in our algorithm.
As $\delta\geq 1$, we assume that $\delta:=(1+\epsilon)$. The key idea is to determine the neighborhood of a vertex in $B_{i,3}$ and then progressively expand this neighborhood until we obtain sets $D_{i,1}^{j}$ and $E_{i,1}^{j}$ (where $E_{i,1}^{j}\supseteq D_{i,1}^{j}$) that satisfy Theorem 5.2. Once this is achieved, we remove the current neighborhood and repeat the process for the remaining graph. Note that $E_{i,1}^{j}$ is not a 2-distance subset, but $D_{i,1}^{j}$ is. The complete procedure is described below (see the pseudocode in Algorithm 1):
  • Initially, set $i\leftarrow 1$.
  • Stage 1: Initialize $j\leftarrow 1$, and set $B_{i,1}^{j}\leftarrow B_{i,1}$, $B_{i,2}^{j}\leftarrow B_{i,2}$, and $B_{i,3}^{j}\leftarrow B_{i,3}$.
  • Stage 2: Choose an arbitrary vertex $v_{i}^{j}$ from $B_{i,1}^{j}$.
  • For $r=0,1,\dots$, consider the $r$-th neighborhood $\mathrm{N}_{i,j}^{r}[v_{i}^{j},B_{i,3}^{j}]$. Starting with $\mathrm{N}_{i,j}^{0}[v_{i}^{j},B_{i,3}^{j}]$ and compute the minimum selective subset while inequality (2) holds.
    $\lvert S^{min}(X_{i,j}^{r+2})\rvert>\delta\cdot\lvert S^{min}(X_{i,j}^{r})\rvert$ (2)
    Here $\mathrm{N}_{i,j}^{r}$ (rather than $\mathrm{N}_{i}^{r}$) denotes the $r$-th neighborhood used to compute $D_{i,1}^{j}$ and $E_{i,1}^{j}$ from $B_{i,3}$. The same convention applies to $X_{i,j}^{r}$ and $Y_{i,j}^{r}$.
  • Let $\overline{r_{i,j}}$ be the smallest $r$ for which inequality (2) is violated, i.e.,
    $\lvert S^{min}(X_{i,j}^{\overline{r_{i,j}}+2})\rvert\leq\delta\cdot\lvert S^{min}(X_{i,j}^{\overline{r_{i,j}}})\rvert$.
  • Update the sets as follows:
    $D_{i,1}^{j}$ $\leftarrow X_{i,j}^{\overline{r_{i,j}}},$
    $E_{i,1}^{j}$ $\leftarrow X_{i,j}^{\overline{r_{i,j}}+2},$
    $B_{i,3}^{j+1}$ $\leftarrow B_{i,3}^{j}\setminus\mathrm{N}_{i,j}^{\overline{r_{i,j}}+2}[v_{i}^{j},B_{i,3}^{j}],$
    $B_{i,1}^{j+1}$ $\leftarrow B_{i,1}^{j}\setminus X_{i,j}^{\overline{r_{i,j}}+2},$
    $B_{i,2}^{j+1}$ $\leftarrow B_{i,2}^{j}\setminus Y_{i,j}^{\overline{r_{i,j}}+2},$
    $j$ $\leftarrow j+1.$
  • Repeat the process from Stage 2 until $B_{i,3}^{j}=\emptyset$.
  • Once $B_{i,3}^{j}$ becomes empty, set $i\leftarrow i+1$ and repeat the process from Stage 1 until $i=k+1$.
Suppose the sets $D_{i,1}^{1},$ $D_{i,1}^{2},$ $\dots,D_{i,1}^{t_{i}}$ and $E_{i,1}^{1},$ $E_{i,1}^{2},$ $\dots,E_{i,1}^{t_{i}}$ are returned from the above algorithm for $1\leq i\leq k$. We establish the following lemmas.
Lemma 8
The sets $\{D_{i,1}^{1},D_{i,1}^{2},\dots,D_{i,1}^{t_{i}}\}$, where $1\leq i\leq k$, obtained from the above algorithm, form a collection of 2-distance subsets.
Lemma 9
For the collection of sets $\{E_{i,1}^{1},\dots,E_{i,1}^{t_{i}}\}$ obtained from the above algorithm, the union $S=\bigcup_{i=1}^{k}\bigcup_{j=1}^{t_{i}}S^{min}(E_{i,1}^{j})$ forms a selective subset of $G$.
Combining Lemmas 8 and 9 with Theorem 5.2, we obtain the following theorem directly:
Theorem 5.3
The above algorithm produces a selective subset $\bigcup_{i=1}^{k}$ $\bigcup_{j=1}^{t_{i}}$ $S^{min}($ $E_{i,1}^{j})$ of size at most $(1+\epsilon)$ times the size of the minimum selective subset of $G$.
Until now, we have considered general graphs rather than unit disk graphs. Thus, Theorem 5.3 holds for general graphs.
5.3 Finding $S^{min}(E_{i,1}^{j})$ on Unit Disk Graphs
The only remaining task is to compute $S^{min}(E_{i,1}^{j})$ in time $n^{f(\epsilon)}$ on unit disk graphs. We assume that $F_{i,j}=\mathrm{N}_{i,j}^{\overline{r_{i,j}}+2}[v_{i}^{j},B_{i,3}^{j}]$. According to the above algorithm,
$F_{i,j}=X_{i,j}^{\overline{r_{i,j}}+2}\cup Y_{i,j}^{\overline{r_{i,j}}+2}=E_{i,1}^{j}\cup Y_{i,j}^{\overline{r_{i,j}}+2}.$
We first show that the size of $S^{min}(E_{i,1}^{j})$ is at most the size of the Maximum Independent Set of $F_{i,j}$. This provides a bound on $E_{i,1}^{j}$. The Maximum Independent Set of a graph $G$ is the largest set of vertices such that no two vertices in the set are adjacent. One might think that $S^{min}(E_{i,1}^{j})$ is the same as a Minimum Dominating Set of $F_{i,j}$. However, by Theorem 5.1, some vertices in $B_{i,3}\setminus B_{i,1}$ may have no adjacent vertex (including themselves) in $S(B_{i,1})$. Therefore, with this bound, it suffices to compute each $S^{min}(E_{i,1}^{j})$.
Lemma 10
The size of $S^{min}(E_{i,1}^{j})$ is at most the size of the maximum independent set of $F_{i,j}$.
Now, we apply the method described in [22] to find a Maximum Independent Set in $F_{i,j}$ for unit disk graphs.
Lemma 11
For any unit disk graph $U$ and an independent set $I^{r}\subseteq\mathrm{N}_{i,j}^{r}[v_{i}^{j},$ $B_{i,3}^{j}]$, we have $\lvert I^{r}\rvert=(2r+1)^{2}=\mathcal{O}(r^{2}).$
Using Lemmas 10 and $1$, we derive the following theorem directly:
Theorem 5.4
The size of the minimum selective subset (local) of $E_{i,1}^{j}$ satisfies
$\lvert S^{\min}(E_{i,1}^{j})\rvert=\mathcal{O}(r^{2}).$
Lemma 12
There exists a constant $d(\delta)$, depending on $\delta=(1+\epsilon)$, such that $\overline{r_{i,j}}\leq d(\delta)$. The running time of our algorithm to compute a PTAS is $\mathcal{O}(n^{d^{2}})$, where $|V(U)|=n$, $d(\epsilon)=\mathcal{O}\Big(\frac{1}{\epsilon^{2}}\log\frac{1}{\epsilon}\Big)$, and $0<\epsilon<\frac{1}{10}$.
6 APX-hardness of MSS on Circle Graphs
The vertex set of a circle graph is a set of chords of a given circle, and if two chords intersect, the corresponding vertices share an edge. We obtain a “gap-preserving” reduction from the MAX-3SAT(8) problem to a circle graph using the Minimum Dominating Set (MDS) problem on circle graphs. The MAX-3SAT(8) problem is as follows [25]:
We are given a set of $n$ variables $\mathcal{X}=\{x_{1},\dots,x_{n}\}$ and $m$ clauses $\mathcal{C}=\{c_{1},\dots,c_{m}\}$ such that each clause has at most $3$ literals and each variable occurs in at most 8 clauses. The objective is to find a truth-assignment of the variables in $\mathcal{X}$ that maximizes the number of clauses in $\mathcal{C}$ satisfied. MAX-3SAT(8) is APX-hard, and the MDS problem on circle graphs is also APX-hard [10].
Consider an instance $\phi$ of MAX-3SAT(8). Let $\text{SAT}(\phi)$ represent the maximum fraction of clauses in $\phi$ that can be satisfied. For a given graph $G$, let $\gamma(G)$ denote the cardinality of its Minimum Dominating Set. The paper [10] reduces an instance of MAX-3SAT(8) to a circle graph $G$ such that $\lvert V(G)\rvert=m+56n+4$ and states the following theorem:
Theorem 6.1
A polynomial-time reduction transforms an instance $\phi$ of MAX-3SAT(8), consisting of $n$ variables and $m$ clauses, into a circle graph $G$ such that
$\text{SAT}(\phi)=1\implies\gamma(G)\leq 16n+2,$
$\text{SAT}(\phi)<\alpha\implies\gamma(G)>16n+2+\frac{(1-\alpha)m}{8},\quad\text{for any }0<\alpha<1.$
We now reduce the graph $G$ into a graph $H$ as follows:
Figure 4: Colors: blue $\equiv$ normal chord, and red $\equiv$ dotted chord. An example of the reduction: each chord of $G$ is colored blue in $H$. Each red chord in $H$ is adjacent only to a chord of blue color.
Reduction. Define $V(H)=V(G)\cup V(X)$ and $E(H)=E(G)\cup E(X)$, where $X$ is a subgraph with vertex set $V(X)$ and edge set $E(X)$ are constructed as follows (see Figure 4). Initially, set $V(X):=\emptyset$ and $E(X):=\emptyset$. Assign color 0 to all chords in $V(G)$, i.e., $C(V(G))=\{0\}$. For each chord $u_{i}\in V(G)$, introduce a corresponding chord $v_{i}\in V(X)$ such that $C(v_{i})=1$. Position $v_{i}$ so that it intersects only $u_{i}$ (meaning $v_{i}$ has degree 1 in $H$). These newly created edges are then included in $E(X)$. As a result, we obtain $\lvert V(H)\rvert=2m+112n+8$, $\lvert V(X)\rvert=m+56n+4$ and $\lvert C(V(H))\rvert=2$.
Lemma 13
$G$ has a dominating set of size $t$ if and only if $H$ has a selective subset of size $m+56n+4+t$.
Let $\lvert S^{min}(H)\rvert$ denote the cardinality of the minimum selective subset of $H$. We now show the “gap-preserving” reduction from MAX-3SAT(8) to the graph $H$ in the following theorem:
Theorem 6.2
The MSS problem is APX-hard in circle graphs.
Remark 3
Our APX-hard result holds not only for circle graphs with two colors; indeed, any color (other than blue) can replace the red chords adjacent to the blue ones.
7 Conclusion
It is still open regarding whether MSS problem is NP-complete on unit disk graphs when $c$ is constant. This raises the open question of whether FPT is possible when the number of colors $c$ is treated as a parameter. Another direction to explore is whether FPT is possible when the number of blocks $k$ is treated as a parameter. Additionally, since MSS is APX-hard on circle graphs, it is natural to ask whether a constant-factor approximation, or a $(2+\epsilon)$-approximation is possible for circle graphs, in contrast to a $(1+\epsilon)$-approximation. FPT for MSS problem with respect to parameters $c$ or $k$ also remains an open question in the context of circle graphs.
Moreover, exploring the complexity of MSS on additional graph classes—such as circular-arc graphs, chordal graphs, and permutation graphs—could uncover new tractable cases or reveal deeper structural insights. These graph families often arise in scheduling, bioinformatics, and network analysis, and understanding MSS within these domains may have practical implications as well.
Acknowledgments
I sincerely thank Bodhayan Roy and Aritra Banik for their support and guidance. Also, the motivation for this work arose from discussions on the MCS paper with Aritra Banik, facilitated by the ACM India Anveshan Setu Fellowship program in 2024. I also thank the anonymous reviewers for their thoughtful comments and suggestions, which helped enhance the quality and presentation of the paper.
Appendix 0.A Proof of the Lemma 1
Proof
Suppose $\mathrm{N}[v,B_{i,3}]\cap S=\emptyset$. Let $v\in V_{\ell}$ for some $1\leq\ell\leq c$. Since each vertex of $B_{i,1}$ must have at least one adjacent vertex of a different color, there must exist a vertex $w\notin B_{i}$ such that $\mathrm{d}(v,w)=1$ and $C(w)\neq C(v)$. If no vertex of $\mathrm{N}[v,B_{i,3}]$ is in $S$, then $v$ has $w$ as its only nearest neighbor in $S\cup(V(G)\setminus V_{\ell})$, and $C(v)\neq C(w)$, contradicting the assumption that $S$ is a selective subset.∎
Appendix 0.B Proof of the Lemma 2
Proof
Let $D$ be a Dominating Set of $G$ of size $s$. Define $S=D\cup\{z\}$; then $\lvert S\rvert=s+1$. We claim that $S$ is a selective subset of $G^{\prime}$. If not, then there exists a vertex $u\in V(G)\setminus S$ such that $\mathrm{d}(u,S)>\mathrm{d}(u,z)$ which implies that $D$ is not a Dominating Set of $G$.
Conversely, suppose $S$ is a selective subset of size $s+1$ in the graph $G^{\prime}$. Due to Lemma 1, $z\in S$. Let $D=S\setminus\{z\}$. We claim that $D$ is a Dominating Set of $G$. If not, then there exists a vertex $u\in V(G)$ such that $\mathrm{N}[u,V(G)]\cap D=\emptyset$, which would imply $\mathrm{d}(u,S)>\mathrm{d}(u,z)$, contradicting the fact that $S$ is a selective subset. Hence our lemma is proved.∎
Appendix 0.C Proof of the Theorem 3.1
Proof
We recall that the Minimum Dominating Set problem is known to be APX-hard on general graphs; that is, there exists a constant $\delta>0$ such that no polynomial-time algorithm can approximate the solution within a factor of $\delta\cdot\log n$ unless $\textup{{P}}=\textup{{NP}}$.
By Lemma 2, there is an approximation-preserving reduction from the Minimum Dominating Set problem to the MSS problem. In particular, given any instance of Minimum Dominating Set, one can construct in polynomial time an instance of MSS such that any approximation algorithm for MSS would translate to an approximation algorithm of comparable quality for Minimum Dominating Set.
Therefore, since Minimum Dominating Set is APX-hard on general graphs, it follows immediately from the reduction that the MSS problem inherits the same hardness of approximation. Hence, the MSS problem is log-APX-hard on general graphs.∎
Appendix 0.D Proof of the Lemma 3
Proof
Let $D$ be a dominating set of $U$ of size $t$. Each vertex in $V(X)$ forms a block, since each vertex has a unique color. By Lemma 1, every selective subset $S$ must include all vertices in $V(X)$, i.e., $V(X)\subseteq S$. Therefore, to construct a selective subset of $U^{\prime}$, we take all vertices in $V(X)$ along with the corresponding vertices of $D$. Let $S=V(X)\cup D$. Since every vertex in $V(U)$ is either in $D$ or adjacent to a vertex in $D$, and since each vertex in $V(X)$ is already in $S$, we conclude that $S$ is a selective subset of size $nm+t$. This holds because for each vertex $v\in V(U^{\prime})\setminus S$, $\mathrm{d}(v,S)=1=\mathrm{d}(v,V(X))$.
Conversely, suppose $S$ is a selective subset of $U^{\prime}$ of size $nm+t$. By Lemma 1, $S$ must contain all vertices of $V(X)$. Define $D^{\prime}=S\setminus V(X)$. We claim that $D^{\prime}$ is a dominating set of $U$ with $\lvert D^{\prime}\rvert=t$. For every vertex $u\in V(U)$, $S$ ensures that either $u\in S$ or $u$ has an adjacent vertex in $S$ of the same color. This holds since $\mathrm{d}(u,V(X))=1$ and $C(u)\notin C(V(X))$. Therefore, the corresponding vertices in $D^{\prime}$ must satisfy the same conditions. Thus, $D^{\prime}$ is a dominating set of $U$ of size tt.∎
Appendix 0.E Proof of the Theorem 4.1
Proof
It is straightforward to see that MSS belongs to NP, since a given solution can be verified in polynomial time. By Lemma 3, we established a reduction from the NP-complete MDS problem in UDG [8]. Hence, MSS is NP-complete on unit disk graphs.∎
Appendix 0.F Proof of the Lemma 4
Proof
Suppose that $v\in B_{i}\setminus B_{i,3}$ and $v\notin S^{min}$. Let $v\in V_{\ell}$ and $w\in\mathrm{\hat{N}}(v,B_{i,1})$. By Lemma 1, we have $\mathrm{N}[w,B_{i,3}]\cap S^{min}\neq\emptyset$. Since $v\notin B_{i,3}$, $\mathrm{N}[w,B_{i,3}]\cap(S^{min}\setminus\{v\})\neq\emptyset$. Therefore for any $j\neq i$, $\mathrm{d}(v,B_{j})\geq\mathrm{d}(v,w)+\mathrm{d}\big(w,\mathrm{N}(w,B_{i,3})\big)$.
This implies that $v$ must have a nearest vertex in $S^{min}\setminus\{v\}\cup(V(G)\setminus V_{\ell})$ with the same color as $C(v)$. Hence, no vertex from $B_{i}\setminus B_{i,3}$ is in $S^{min}$.
Since no vertex $v\in B_{i}\setminus B_{i,3}$ is in $S^{min}$, it follows that $S^{min}\subseteq\bigcup_{i=1}^{c}B_{i,3}$.∎
Appendix 0.G Proof of the Theorem 5.1
Proof
Let $C(B_{i})=\ell$. By Lemma 4, every minimum selective subset is contained in $\bigcup_{i=1}^{k}B_{i,3}$. Hence, it remains to show that every vertex in $B_{i,3}\setminus S(B_{i,1})$ has a nearest neighbor of its own color in $S(B_{i,1})\cup(V(G)\setminus V_{\ell})$.
Consider a vertex $w\in B_{i,3}\setminus S(B_{i,1})$. We distinguish two cases:
Case 1. $w\in B_{i,1}$. By Lemma 1, at least one adjacent vertex of $w$ in $B_{i,3}$ must belong to $S(B_{i,1})$. So, $w$ has a nearest neighbor in $S(B_{i,1})\cup(V(G)\setminus V_{\ell})$ of its own color.
Case 2. $w\in B_{i,2}$. Let $u\in B_{i,1}$ be an adjacent vertex of $w$. Since $w\notin S(B_{i,1})$ and $u\in B_{i,1}$, by Lemma 1, it follows that $(\mathrm{N}[u,B_{i,3}]\setminus\{w\})\cap S(B_{i,1})\neq\emptyset$. Because $\mathrm{d}(w,u)=1$, we have $\mathrm{d}(w,B_{j})\;\geq\;\mathrm{d}(w,u)+\mathrm{d}\big(u,(\mathrm{N}(u,B_{i,3})\setminus\{w\})\big)$ for all $j\neq i$. Thus, $w$ has a nearest neighbor of its own color in $(\mathrm{N}[u,B_{i,3}]\setminus\{w\})\cap S(B_{i,1})$.
Therefore, $S(B_{i,1})$ is also a selective subset for $B_{i}$. Consequently, constructing $S=\bigcup_{i=1}^{k}S(B_{i,1})$ with size at most a $(1+\epsilon)$-factor of the global optimum is sufficient.∎
Appendix 0.H Proof of the Lemma 5
Proof
Since $\mathrm{d}(D_{i,1}^{j},D_{i,1}^{l})>2$, we have $\mathrm{d}(\mathrm{N}[D_{i,1}^{j},$ $B_{i,3}],$ $\mathrm{N}[D_{i,1}^{l},$ $B_{i,3}])\geq 1$, which implies that their neighborhoods are disjoint. Hence, the first claim holds.
Since $S^{min}(D_{i,1}^{j})$ $\subseteq\mathrm{N}[D_{i,1}^{j},B_{i,3}]$ and $S^{min}(D_{i,1}^{l})$ $\subseteq\mathrm{N}[D_{i,1}^{l},$ $B_{i,3}]$, the second claim follows immediately.
Finally, since $S^{min}(D_{i,1}^{j})\cap S^{min}(D_{i,1}^{l})=\emptyset$, the third claim also holds. ∎
Appendix 0.I Proof of the Lemma 6
Proof
From Lemma 4, every vertex in $S^{\min}$ lies within the neighborhood $\bigcup_{i=1}^{k}\mathrm{N}[B_{i,1},B_{i,3}]$. Lemma 1 further states that for each vertex $v\in B_{i,1}$, either $v\in S^{\min}$, or at least one of its adjacent vertices $u\in B_{i,3}$ must be included in $S^{\min}$.
Moreover, by the proof of Theorem 5.1, for any vertex $x\in\bigcup_{i=1}^{k}\mathrm{N}[B_{i,1},B_{i,3}]\setminus S^{\min}$, if $x\in V_{\ell}$, then there exists a nearest vertex $y\in S^{\min}\cup(V(G)\setminus V_{\ell})$ such that $C(x)=C(y)$. Therefore, $x$ satisfies the definition of the selective subset problem.
Therefore, for every 2-distance subset $D_{i,1}^{j}$, the set $S^{\min}\cap\mathrm{N}[D_{i,1}^{j},B_{i,3}]$ forms a local selective subset for $D_{i,1}^{j}$, since every vertex $v\in D_{i,1}^{j}\subseteq B_{i,1}$ is either in the set $S^{\min}\cap\mathrm{N}[D_{i,1}^{j},B_{i,3}]$ or has at least one adjacent vertex in $\mathrm{N}[D_{i,1}^{j},B_{i,3}]$.∎
Appendix 0.J Proof of the Lemma 7
Proof
By Lemma 5, since the neighborhoods of any two 2-distance subsets are disjoint, the corresponding local minimum selective subsets are disjoint. Therefore, each $S^{min}(D_{i,1}^{j})$ contributes uniquely to $S^{min}$.
Moreover, by Lemma 6, $S^{min}\cap\mathrm{N}[D_{i,1}^{j},B_{i,3}]$ must be a local selective subset of $D_{i,1}^{j}$. Also, $S^{min}(D_{i,1}^{j})$ is the minimum selective subset for $D_{i,1}^{j}$. Therefore,
$\lvert S^{min}(D_{i,1}^{j})\rvert\leq\lvert S^{min}\cap\mathrm{N}[D_{i,1}^{j},B_{i,3}]\rvert$.
Summing over all $i$ and $j$, we get:
$\sum_{i=1}^{k}\sum_{j=1}^{t_{i}}\lvert S^{min}(D_{i,1}^{j})\rvert\leq\sum_{i=1}^{k}\sum_{j=1}^{t_{i}}\lvert S^{min}\cap\mathrm{N}[D_{i,1}^{j},B_{i,3}]\rvert\leq\lvert S^{min}\rvert$.
Appendix 0.K Algorithm 1
Input: Graph $G=(V(G),E(G))$, initial sets $B_{i,1},B_{i,2},B_{i,3}$ for $1\leq i\leq k$, and parameter $\delta$.
Output: Sets $D_{i,1}^{j}$ and $E_{i,1}^{j}$ for all $i,j$.
1
2Initialize $i\leftarrow 1$;
3
4while $i\leq k$ do
5   Initialize $j\leftarrow 1$;
6   Set $B_{i,1}^{j}\leftarrow B_{i,1}$, $B_{i,2}^{j}\leftarrow B_{i,2}$, $B_{i,3}^{j}\leftarrow B_{i,3}$;
7
8 while $B_{i,3}^{j}\neq\emptyset$ do
9      Select an arbitrary vertex $v_{i}^{j}\in B_{i,1}^{j}$;
10
11    for $r=0,1,2,\dots$ do
12         Compute $\mathrm{N}_{i,j}^{r}[v_{i}^{j},B_{i,3}^{j}]$;
13         Compute $S^{min}(X_{i,j}^{r+2})$;
14
15       if $\lvert S^{min}(X_{i,j}^{r+2})\rvert\leq\delta\cdot\lvert S^{min}(X_{i,j}^{r})\rvert$ then
16            Break;
17
18         end if
19
20      end for
21
22     Set $\overline{r_{i,j}}$ as the smallest $r$ for which the condition is violated;
23
24     Update sets:;
25    $D_{i,1}^{j}\leftarrow X_{i,j}^{\overline{r_{i,j}}}$;
26    $E_{i,1}^{j}\leftarrow X_{i,j}^{\overline{r_{i,j}}+2}$;
27    return Selective subset for $D_{i,j},E_{i,j}$ $B_{i,3}^{j+1}\leftarrow B_{i,3}^{j}\setminus\mathrm{N}_{i,j}^{\overline{r_{i,j}}+2}[v_{i}^{j},B_{i,3}^{j}]$;
28    $B_{i,1}^{j+1}\leftarrow B_{i,1}^{j}\setminus X_{i,j}^{\overline{r_{i,j}}+2}$;
29    $B_{i,2}^{j+1}\leftarrow B_{i,2}^{j}\setminus Y_{i,j}^{\overline{r_{i,j}}+2}$;
30
31     Increment $j\leftarrow j+1$;
32
33   end while
34
35  Increment $i\leftarrow i+1$;
36
37 end while
Algorithm 1 Algorithm for Computing $D_{i,1}^{j}$ and $E_{i,1}^{j}$
Appendix 0.L Proof of the Lemma 8
Proof
We prove this by mathematical induction on $j=1,\dots,t_{i}$ for each $i$.
For the base case $j=1$, we have
$B_{i,3}^{2}=B_{i,3}^{1}\setminus\mathrm{N}_{i,1}^{\overline{r_{i,j}}+2}[v_{i}^{1},B_{i,3}^{1}].$
Since
$\mathrm{N}[\mathrm{N}[D_{i,1}^{1},B_{i,3}^{1}]]=\mathrm{N}[\mathrm{N}[\mathrm{N}_{i,1}^{\overline{r_{i,j}}}[v_{i}^{1},B_{i,3}^{1}]]]=\mathrm{N}_{i,1}^{\overline{r_{i,j}}+2}[v_{i}^{1},B_{i,3}^{1}],$
it follows that $\{D_{i,1}^{1},B_{i,1}^{2}\}$ form a 2-distance subset because $\mathrm{d}(D_{i,1}^{1},B_{i,1}^{2})>2$ and $D_{i,1}^{1},B_{i,1}^{2}\subseteq B_{i,1}^{1}$.
Now, as the induction hypothesis, assume that $\{D_{i,1}^{1}$ $\dots,$ $D_{i,1}^{j-1},$ $B_{i,1}^{j}\}$ forms a 2-distance subset. We need to show that $\{D_{i,1}^{1},\dots,D_{i,1}^{j},B_{i,1}^{j+1}\}$ is also a 2-distance subset.
Since
$B_{i,3}^{j+1}=B_{i,3}^{j}\setminus\mathrm{N}_{i,1}^{\overline{r_{i,j}}+2}[v_{i}^{j},B_{i,3}^{j}]=B_{i,3}^{j}\setminus\mathrm{N}[\mathrm{N}[D_{i,1}^{j},B_{i,3}^{j}]],$
this completes the proof.∎
Appendix 0.M Proof of the Theorem 5.2
Proof
By summing over all indices, we obtain:
[…]
where the final step follows from Lemma 7.∎
Appendix 0.N Proof of the Lemma 9
Proof
Since $B_{i,1}^{j+1}=B_{i,1}^{j}\setminus E_{i,1}^{j}$, where $E_{i,1}^{j}=X_{i,j}^{\overline{r_{i,j}}+2}$ and $E_{i,1}^{j}\subset B_{i,1}^{j}$, we have $B_{i,1}^{j}=B_{i,1}^{j+1}\cup E_{i,1}^{j}.$ This implies that $\bigcup_{j=1}^{t_{i}}E_{i,1}^{j}=B_{i,1}$.
Thus, $\bigcup_{j=1}^{t_{i}}S^{min}(E_{i,1}^{j})$ is a selective subset of $B_{i,1}$ for $1\leq i\leq k$. Consequently, the set
$\bigcup_{i=1}^{k}\bigcup_{j=1}^{t_{i}}S^{min}(E_{i,1}^{j})$
forms a selective subset of GG.∎
Appendix 0.O Proof of the Lemma 10
Proof
According to Lemma 1, for each vertex $v\in E_{i,1}^{j}\subseteq F_{i,j}$, either $v\in S^{min}(E_{i,1}^{j})$ or at least one of its neighbors in $F_{i,j}$ must be in $S^{min}(E_{i,1}^{j})$. However, vertices in $F_{i,j}\setminus S^{min}(E_{i,1}^{j})$ may not have any adjacent vertex in $S^{min}(E_{i,1}^{j})$, as shown in the proof of Theorem 5.1. Therefore, $S^{min}(E_{i,1}^{j})$ may not itself be a dominating set of $F_{i,j}$, but its size is at most that of a minimum dominating set of $F_{i,j}$.
Moreover, by the proof of Theorem 5.1, if $u\in F_{i,j}\setminus S^{min}(E_{i,1}^{j})$ and $C(u)=\ell$, then $u$ has a nearest neighbor of the same color in $S^{min}(E_{i,1}^{j})\cup(V(G)\setminus V_{\ell})$. That is, every vertex in $F_{i,j}\setminus S^{min}(E_{i,1}^{j})$ satisfies the condition of a selective subset.
Here, $\gamma(F_{i,j})$ denotes the size of a minimum dominating set of $F_{i,j}$, and $\alpha(F_{i,j})$ denotes the size of a maximum independent set of $F_{i,j}$.
Thus, we have
$|S^{min}(E_{i,1}^{j})|\leq\gamma(F_{i,j}).$
Since every maximum independent set in $F_{i,j}$ is a dominating set, it follows that
$\gamma(F_{i,j})\leq\alpha(F_{i,j}).$
Consequently,
$|S^{min}(E_{i,1}^{j})|\leq\alpha(F_{i,j}),$
as claimed.
Appendix 0.P Proof of the Lemma 11
Proof
By the definition of unit disk graphs, we have:
$u,v\in\mathrm{N}_{i,j}^{r}[v_{i}^{j},B_{i,3}^{j}]\quad\iff\quad\|f(u)-f(v)\|\leq 2r.$ (3)
Thus, $I^{r}$ consists of mutually disjoint unit disks within a disk of radius $2r+1$ centered at $f(v)$.
Since the maximum number of such disjoint unit disks that can fit in a disk of radius $2r+1$ is given by:
$\lvert I^{r}\rvert\leq\frac{\pi(2r+1)^{2}}{\pi}=(2r+1)^{2}=\mathcal{O}(r^{2}),$
the result follows.∎
Appendix 0.Q Proof of the Lemma 12
Proof
For any $r<\overline{r_{i,j}}$, condition (2) implies the following inequalities depending on whether $r$ is even or odd:
When $r$ is even:
$\begin{split}(2r+1)^{2}\geq\lvert S^{min}(X_{i,j}^{r+2})\rvert>\delta\lvert S^{min}(X_{i,j}^{r})\rvert\\ >\dots>\delta^{\frac{r}{2}}\lvert S^{min}(X_{i,j}^{0})\rvert=\delta^{\frac{r}{2}}.\end{split}$ (4)
When $r$ is odd:
$(2r+1)^{2}$ $\geq\lvert S^{min}(X_{i,j}^{r+2})\rvert>\delta\lvert S^{min}(X_{i,j}^{r})\rvert$ (5)
$>\dots>\delta^{\frac{r}{2}}\lvert S^{min}(X_{i,j}^{1})\rvert=\delta^{\frac{r}{2}}.$
Since $\delta>1$, the above inequalities hold until condition (2) becomes violated. Therefore, when $r<\overline{r_{i,j}}$ the above inequalities are violated (that is, $(2\overline{r_{i,j}}+1)^{2}<\delta^{\frac{\overline{r_{i,j}}}{2}}$), and this first violation determines the value of $\overline{r_{i,j}}$, which depends only on $\delta$, not on the size of $V(U)$. To prove this, we now show that $(2d+1)^{2}<(1+\epsilon)^{d}$ where $d=\mathcal{O}\Big(\frac{1}{\epsilon^{2}}\log\frac{1}{\epsilon}\Big)$. Note that we use $d$ instead of $\overline{r_{i,j}}$ for simplicity of calculation.
For any $d\geq 1$, we first prove that
$(2d+1)^{2}<(3d)^{2}<(1+\epsilon)^{d}.$
The first inequality holds for all $d\geq 1$. To prove the second inequality, it suffices to show:
$2\log(3d)<d\log(1+\epsilon).$
This implies:
$\frac{2}{d}\log(3d)+\epsilon^{2}<\log(1+\epsilon)+\epsilon^{2}.$
Substituting $d=\frac{1}{\epsilon^{2}}\log\frac{1}{\epsilon}$ into $\frac{2}{d}\log(3d)+\epsilon^{2}$, we obtain:
$\frac{2}{d}(\log 3+\log d)+\epsilon^{2}$ $=\frac{2\epsilon^{2}}{\log\frac{1}{\epsilon}}\left(\log 3+\log\frac{1}{\epsilon^{2}}+\log\log\frac{1}{\epsilon}\right)+\epsilon^{2}$ (6)
$=2\epsilon^{2}\left(\frac{\log 3}{\log\frac{1}{\epsilon}}+2\frac{\log\frac{1}{\epsilon}}{\log\frac{1}{\epsilon}}+\frac{\log\log\frac{1}{\epsilon}}{\log\frac{1}{\epsilon}}+\frac{1}{2}\right)$
[…]
Since $0<\epsilon<\frac{1}{10}$, it follows that:
$2\epsilon^{2}(\log 3+3.5)<10\epsilon^{2}<\epsilon,$ (7)
As established in [15], for $0\leq\epsilon<\frac{1}{2}$, the following inequality holds
$\epsilon\leq\log(1+\epsilon)+\epsilon^{2}$ (8)
Substituting inequalities (7) and (8) into inequality (6), we obtain
$\frac{2}{d}\log(3d)+\epsilon^{2}<\log(1+\epsilon)+\epsilon^{2}.$
Moreover, by Theorem 5.4, $S^{min}(E_{i,1}^{j})$ can be computed in $\mathcal{O}(n^{d^{2}})$, where $d(\epsilon)=\mathcal{O}(\frac{1}{\epsilon^{2}}\log\frac{1}{\epsilon})$ and $0<\epsilon<\frac{1}{10}$.∎
Appendix 0.R Proof of the Lemma 13
Proof
If $D$ is a dominating set of $G$ of size $t$, then $S=D\cup V(X)$ forms a selective subset of $H$. This holds because each chord in $V(X)$ forms a block, and by Lemma 1, we have $V(X)\subset S$. Moreover, for every vertex $v\in V(H)\setminus S$, we have $\mathrm{d}(v,S)=\mathrm{d}(v,V(X))=1$.
Conversely, if $S$ is a selective subset of size $m+56n+4+t$, then by Lemma 1, we have $V(X)\subset S$, which implies that $\lvert S\setminus V(X)\rvert=t$. The set $S\setminus V(X)$ forms a dominating set of $G$, since for each vertex $v\in V(G)$, we have $\mathrm{d}(v,V(X))=1$ in $H$. Therefore, either $v\in S\setminus V(X)$ or at least one neighbor of $v$ from $V(G)$ must be included in $S\setminus V(X)$. This completes the proof.∎
Appendix 0.S Proof of the Theorem 6.2
Proof
Using Lemma 13, we have $S^{min}(H)=m+56n+4+\gamma(G)$. We use this equation in Theorem 6.1 for any $0<\alpha<1$, and we get
$\text{SAT}(\phi)=1\implies\lvert S^{min}(H)\rvert\leq 16n+2+(m+56n+4),$
$\text{SAT}(\phi)<\alpha\implies\lvert S^{min}(H)\rvert>16n+2+(m+56n+4)+\frac{(1-\alpha)m}{8}.$
This clearly indicates that the gap-preserving reduction works. Therefore, this reduction shows that MSS is APX-hard in circle graphs. Hence, the selective subset problem does not admit a PTAS unless $\textup{{P}}=\textup{{NP}}$.∎